OpenStudy (anonymous):

laplace transform of cos t+ cos 2t from Pi

OpenStudy (anonymous):

What are we doing here exactly?

OpenStudy (aravindg):

cos 2t =2cos^2 t-1

OpenStudy (aravindg):

you will get a quadratic in t

OpenStudy (aravindg):

solve for t

OpenStudy (anonymous):

Huueefetcha!!

OpenStudy (aravindg):

note that Pi<t<2Pi

OpenStudy (anonymous):

how will the pi range affects the ans if i simply use formula to take L.T of the function?

OpenStudy (aravindg):

for example if i get t=5pi/4 i need to convert it into an angle btw pi and -pi

OpenStudy (aravindg):

where cos has same value

OpenStudy (aravindg):

gt it?

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

so i express cos 2t in t first?

OpenStudy (anonymous):

u know the formula \[\text{L{cos (at)}}=\frac{s}{a^2+s^2}\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so without consdering the pi range i get s/(s^2+1)+s/(s^2+4)

OpenStudy (anonymous):

den how do i continue?

OpenStudy (anonymous):

I think you need the shifted step function in there, to truncate your cosines... been a while since I did this..

OpenStudy (anonymous):

how to do that? need to plot out the graph and see?

OpenStudy (anonymous):

checking...

OpenStudy (anonymous):

Laplace transform is defined as F(s)=L[f(t)] = ∫ e^(-st) f(t) dt Let f(t)=g(t)+h(t) where g(t)=cos t and h(t)=cos 2t The integral becomes H(s)=L[h(t)]=∫ e^(-st) cos(2t) dt integrate by parts with u=cos(2t) → du= -2sin(2t) dv=e^(-st) → v= -(1/s)e^(-st) ∫ cos(2t)e^(-st) dt = -(1/s)e^(-st)cos(2t) - (b/s) ∫ e^(-st)sin(2t) dt to solve the integrals on right side integrate by parts with u=sin(2t) → du= 2cos(2t) dv=e^(-st) → v= -(1/s)e^(-st) ∫ cos(2t)e^(-st) dt = -(1/s)e^(-st)cos(2t) - (2/s) [ -(1/s)e^(-st)sin(2t) + (2/s)∫ e^(-st)cos(2t) dt <=> ∫ cos(2t)e^(-st) dt = -(1/s)e^(-st)cos(2t) + (2/s²)e^(-st)sin(2t) - (4/s²)∫ e^(-st)cos(2t) dt solve for ∫ cos(2t)e^(-st) dt (1 + (4/s²)) ∫ cos(2t)e^(-st) dt = -(1/s)e^(-st)cos(2t) + (4/s²)e^(-st)sin(2t) <=> ((s² + 4)/s²) ∫ cos(2t)e^(-st) dt = -(1/s)e^(-st)cos(2t) + (2/s²)e^(-st)sin(2t) <=> ∫ cos(2t)e^(-st) dt = [ s²/(s² + 4)]( -(1/s)e^(-st)cos(2t) + (2/s²)e^(-st)sin(2t)) = (-scos(2t) + 2sin(2t))e^(-st)/(s² + 4) Now apply the limits to find L[h(t)] \[L [ \cos(2t) ] = \int\limits_{\pi}^{2\pi} \cos(2t) e^{-st} dt\] Similarly, you can find L[g(t)] = (-scos t+sin t) e^(-st)/(s²+1) Find F(s) by adding the two solutions. L[f(t)]=L[g(t)]+L[h(t)]